OpenStudy (anonymous):

Pls help me with this: Find the exact value cos[2cos^-1(-1/2)]

zepdrix (zepdrix):

\[\large \cos\left[2\color{orangered}{\arccos\left(-\frac{1}{2}\right)}\right]\] Work from the inside out. So what value does this arccosine give us? \[\large \arccos\left(-\frac{1}{2}\right)=\theta \qquad\rightarrow\qquad \cos \theta=-\frac{1}{2}\] Do you know what value of theta this gives us? :o

OpenStudy (anonymous):

@zepdrix yes : 2pi/3 .... so what is the answer?

zepdrix (zepdrix):

\[\large \cos\left[2\color{orangered}{\frac{2\pi}{3}}\right]\] Ok good, you figured out the orange part. Understand how to solve it from here? :) You need to remember your special angles.

OpenStudy (anonymous):

@zepdrix : so the answer is 4pi/3?

zepdrix (zepdrix):

No. The inside simplifies to 4pi/3. What does the cosine of 4pi/3 give you? :o

OpenStudy (anonymous):

@zepdrix : -1/2

zepdrix (zepdrix):

Yay good job \c:/

OpenStudy (anonymous):

Oh I see! Thanks so much!